OpenStudy (anonymous):

a runner starts from rest and speeds up with constant acceleration. if she has gone a distance of 30 m at the point when she reaches a speed of 8m/s what is her acceleration??

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

\[v_f^2=v_i^2+2ad\] (8 m/s)² = 0 + 2a(30 m)